a) If Maria’s utility function is U = A Xayb what is his marginal utility functions for X and Y? Find the MRS between x and y.
1
Expert's answer
2020-07-08T17:48:20-0400
"MU_x=\frac {\delta U}{\delta x}=Aax^{a-1}y^b"


"MU_y=\frac {\delta U}{\delta y}=Abx^{a}y^{b-1}"


"MRS_{x,y}=\frac {MU_x}{MU_y}=\frac {y}{x}"